OpenStudy (anonymous):

How can I solve integral of 2sin(t/2)dt?

OpenStudy (aravindg):

put u=t/2 du=1/2 dt the integral becomes integral 2 sin u 1/2 du = integral sin u du

OpenStudy (anonymous):

Are you looking for \[\int\limits_{} 2\sin(t/2) dt\]?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

This one is easy, really. Can you figure out the antiderivative? The antiderivative of sin is what? And the antiderivative then of t/2 is what? You would be using the chain rule to differentiate to get this function.

OpenStudy (anonymous):

I just solved, and it is -4cos(t/2)+c. But I used u substitution.

OpenStudy (aravindg):

Where did 4 come from ?

OpenStudy (anonymous):

when ask my question before, you forget to multiply du by 2. I think you divided by 2 instead of multiply.

OpenStudy (aravindg):

ya u r right ..sorry for that

OpenStudy (anonymous):

ok :)